# Can Helicity Modulus Be Defined For Boundary Conditions With Finite Twist?

### v1

@@ Abstract
We study the response of a two-dimensional classical XY model to a finite (non-infinitesimal) twist of the boundary conditions. We use Monte Carlo simulations to evaluate the free energy difference between periodic and twisted-periodic boundary conditions and find deviations from the expected quadratic dependence on the twist angle. Consequently, the helicity modulus (spin-stiffness) shows a non-trivial dependence on the twist angle. We show that the deviation from the expected behavior arises because of a degeneracy due to the chirality of the spin-waves which leads to an additional entropy contribution in the quasi-long-range ordered phase. We give an improved prescription for the numerical evaluation of the helicity modulus and resolve some open questions related to anti-periodic boundary conditions. We also discuss applications to discrete spin systems and some experimental scenarios where boundary conditions with finite twist are necessary.
1 Introduction
@@
The study of magnetic systems with the help of spin models such as the Ising or Heisenberg models has resulted in tremendous progress in condensed matter physics. For ferromagnetic systems in the thermodynamic limit, surface energies are negligible, and the choice of boundary conditions can be considered unimportant. On the other hand, in disordered systems such as spin-glasses, effects of various boundary conditions have been a topic of active interest [1, 2, 3]. In the presence of long-range interactions in an Ising spin system, different choices of boundary condition lead to seemingly different thermodynamic behaviors. By choosing artificial coupling-dependent boundary condition, an uncountable number of exotic spin states of the ground state can be generated at any temperature, whereas free boundary conditions, which are considered physical, do not generate the same effect [4]. In the XY or Heisenberg spin glasses, small rotations at the boundary are expected to yield non-smooth change in ground state [5]. However, to the best of our knowledge, a detailed investigation of the boundary dependence in continuous spin systems, similar to the above studies of Ising models, is still missing.
@@
In this paper, we consider the seemingly simpler but surprisingly complex effects of a finite (non-infinitesimal) twist of the boundary conditions on a ferromagnetic XY system. Arguments put forward by Brown and Ciftan [6] in the context of the three-dimensional Heisenberg model suggest that the mixing of states with different chirality plays an important role and affects the observed helicity modulus. However, a quantitative understanding of this effect and its origins has not been achieved yet. We therefore study this question in detail for the two-dimensional classical XY model by means of large-scale Monte Carlo simulations. We find that the free energy cost of a non-infinitesimal twist in the boundary condition in the quasi long-range ordered (QLRO) low-temperature phase deviates from the expected quadratic dependence on the twist angle. Mixing of states with opposite chirality provides an extra entropy contribution which takes the value $\ln(2)$ (in units in which $k_{B}=1$) for a $\pi-$twist. Beyond their intrinsic interest, our results potentially apply to experiments aimed at detecting the Berezinskii–Kosterlitz–Thouless (BKT) transition [7, 8]. They are also important for systems with discrete $Z(N)$ (clock) order parameter symmetry where any twist in the boundary conditions is necessarily non-infinitesimal. In addition, our findings enable novel numerical algorithms for computing the helicity modulus in simulations.
@@
Our paper is organized as follows. In section 2, we introduce the Hamiltonian under various boundary conditions, and we discuss the boundary condition dependence of the free energy. Section 3 contains the details of numerical simulations. In section 4, we present our results for the helicity modulus evaluated for a finite twist in the boundary condition, and we perform a detailed analysis of the twist dependence of thermodynamic observables. We conclude in section 5.
2 The Model
@@
Here, $J>0$ denotes the ferromagnetic exchange interaction (which will be set to unity in the simulations), the sum is over pairs of nearest neighbors on a $d$-dimensional hypercubic lattice, and $\mathbf{S}=(S_{x},S_{y})$ is a two-component unit vector. Equivalently, the XY spins can be represented by their phases, $\phi$ defined via $S_{x}=\cos\phi,~{}S_{y}=\sin\phi$. Twisted boundary conditions can be implemented by fixing the spins at two opposite boundaries at specific orientations (phases) with a fixed angle $\Theta$ between them. Alternatively, we consider the Hamiltonian (1) with periodic boundary conditions and modify the interactions across one of the boundaries to introduce the twist. This is achieved by replacing the interaction terms across the chosen boundary by $-J\cos(\phi_{i}-\phi_{j}-\Theta)$. We call these boundary conditions twisted-periodic, and the usual periodic boundary conditions are recovered for $\Theta=0$.
@@
In a long-range ordered or quasi long-range ordered phase, a twist in the boundary conditions increases the system’s free energy. The system can lower its free energy by distributing the total twist (angular difference) $\Theta$ over the entire sample, i.e., by gradually changing the average orientation of the spins in the bulk. For a system of linear size $L$, the lowest free energy is expected when the average phase changes by $\Theta/L$ between neighboring sites (in the direction the twist is applied). For large $L$, this local phase change is small, and free energy can be expanded in powers of $\Theta/L$. As the free energy difference $\Delta F$ between the twisted and untwisted systems must be an even function of $\Theta$, it takes the form
@@
$\rho_{s}(\Theta)=\frac{2\Delta F}{\Theta^{2}}L^{2-d}.$ (3)
In the limit of an infinitesimal twist, the helicity modulus can be rewritten as $\rho_{s}(\Theta\rightarrow 0)=\left({\partial^{2}F}/{\partial\Theta^{2}}\right)_{\Theta=0}L^{2-d}$. By treating the twist angle $\Theta$ as a parameter in the Hamiltonian, the partial derivative $\left({\partial^{2}F}/{\partial\Theta^{2}}\right)_{\Theta=0}$ can be expressed in terms of appropriate thermodynamic averages in the untwisted system. This leads to the formula [10]
$\left(\frac{\partial^{2}F}{\partial\Theta^{2}}\right)_{\Theta=0}=$ $\frac{1}{L^{2-d}}\sum_{<ij>}J\left\langle\cos(\phi_{i}-\phi_{j})\right\rangle_{\Theta=0}(x_{i}-x_{j})^{2}$ (4)
$-\frac{\beta}{L^{2-d}}\left\langle\left\{\sum_{<ij>}J~{}\sin(\phi_{i}-\phi_{j})(x_{i}-x_{j})\right\}^{2}\right\rangle_{\Theta=0}.$ (5)
where $x_{i}$ is the coordinate of site $i$ in the twisted direction.
As the paramagnetic phase is insensitive to the boundary conditions, the free energy difference decays exponentially with system size, $\Delta F\sim e^{-L/\xi}$ , where $\xi$ is the correlation length. This implies $\rho_{s}=0$ in the paramagnetic phase in the thermodynamic limit. In contrast, $\Delta F$ is expected to scale as $L^{d-2}$ in an ordered or quasi long-range ordered phase, and the helicity modulus is finite. In two dimensions, $\rho_{s}$ is known to have an universal jump at the BKT phase transition. For $d>2$, $\rho_{s}$ vanishes continuously, governed by critical behavior of the phase transition. In the rest of the paper, we focus on two dimensions, but we will comment on higher dimensions in the concluding section.
3 Numerical Simulations
We perform large-scale Monte Carlo simulations to evaluate free energy difference between systems with periodic and twisted-periodic boundary conditions. The free energy cannot be measured directly in a standard Monte Carlo simulation. Instead, it can be evaluated explicitly by integrating the internal energy $U=\langle H\rangle$ over the inverse temperature $\beta=1/T$,
@@ 3 Numerical Simulations
Our simulations employ both the single spin-flip Metropolis algorithm [13, 14] and the Wolff cluster-flip algorithm [15]. For systems with periodic boundary conditions, the efficient Wolff algorithm greatly reduces critical slowing down. Thus, a full MC sweep consists of one Metropolis sweep followed by one Wolff sweep for the case of periodic boundary conditions. However, for twisted-periodic boundary conditions with twist angle $0<\Theta<\pi$, the Wolff algorithm cannot be employed. This stems from the fact that the angle between two spins of a Wolff cluster is not preserved (but rather changes sign) when the cluster is flipped. Consequently, the cluster flip changes the energy of a twisted bond inside the cluster, invalidating the algorithm. In the case of twisted-periodic boundary conditions with $0<\Theta<\pi$, we therefore only employ Metropolis sweeps. For a twist angle of exactly $\pi$, the twisted bond effectively become antiferromagnetic as $\cos(\Delta\phi-\pi)=-\cos(\Delta\phi)$. The energy of an antiferromagnetic bond is invariant under a sign change of $\Delta\phi$, and the Wolff algorithm can be used.
@@
For periodic boundary conditions, we perform up to $2000$ full equilibration sweeps (each consisting of a Metropolis sweep followed by a Wolff sweep) and up to $8000$ full measurement sweeps at each temperature step. For twisted-periodic boundary conditions, we perform up to about 80 0008000080\,00080 000 (Metropolis only) equilibration sweeps and up to about 160 000160000160\,000160 000 measurement sweeps. (As usual, the quality of the equilibration was confirmed by comparing the results of hot and cold starts.) We simulate systems of linear sizes up to $L^{2}=80^{2}$ and average the results over about $5000$ samples. Each sample is subjected to periodic and twisted-periodic boundary conditions, and the resulting free energies are compared to evaluate the helicity modulus (3).
@@
We use an ensemble method (see, e.g., Ref. [17]) to estimate the error of the free energy. We generate a large ensemble of synthetic internal energy curves ${U_{i}(T)=U(T)+r(T)~{}\Delta U(T)}$, where $\Delta U$ is the statistical error obtained from Monte Carlo and $r(T)$ is a random number chosen from a normal distribution of unit variance. Integrating these curves via (6) generates an ensemble of free energies $F_{i}$. Mean and standard deviation of this ensemble are then propagated through (3) to find the helicity modulus and the associated error.
4 Results
@@
We now turn to our results for the response of the two-dimensional XY model to various twists in the boundary conditions. Figure 1 shows the helicity modulus $\rho_{s}$ as a function of temperature for three different cases, viz., an infinitesimal twist, a twist of $\Theta=\pi/6$, and a twist of $\Theta=\pi$.
@@
Figure 1: Helicity modulus $\rho_{s}$ as a function of temperature $T$ for different system sizes $L$. The filled symbols show $\rho_{s}$ for a twist of $\Theta=\pi/6$, obtained from the definition (3) using the free energies (6). The dotted lines denote $\rho_{s}$ for an infinitesimal twist measured in the untwisted system via eq. (5). The helicity modulus values for these two cases agree within their statistical errors (which are comparable to or smaller than the symbol sizes). The straight solid line corresponds to $\rho_{s}=2T/\pi$, it intersects the stiffness curves at temperature $T^{*}(L)$. The critical temperature $T_{c}$ is obtained by extrapolating $T^{*}(L)$ to the thermodynamic limit according to eq. (7). This extrapolation gives $T_{c}=0.891(7)$. In contrast, the helicity modulus values for a twist of $\Theta=\pi$, shown as dash-dotted line for $L=40$, are significantly lower than the other data.
@@
For the infinitesimal twist, $\rho_{s}$ is measured in the untwisted system via eq. (5). For the finite twists, $\rho_{s}$ is obtained from the free energy difference between simulations with twisted and untwisted boundary conditions, as explained in Sec. 3. The resulting $\rho_{s}$ values for the infinitesimal twist and $\Theta=\pi/6$ agree within their statistical errors, giving us additional confidence in our numerical approach.
@@
where $A,b$ are non-universal fitting parameters. We find $T_{c}=0.891(7)$ from twisted boundary conditions with $\Theta=\pi/6$, which agrees with high-accuracy results in the literature [19].
We also studied a twist of $\Theta=\pi$. Unexpectedly, the resulting helicity modulus values are significantly below the values for smaller $\Theta$, as shown in Fig. 1. It is worth emphasizing that this happens even though the layer-to-layer twist $\Theta/L$ remains small compared to unity, justifying the expansion that leads to eq. (2). What is the reason for this surprising discrepancy? Arguments put forward in Ref. [6] suggest that the chirality of the twist plays an important role.
So far we considered thermodynamic states that fulfill the global twist $\Theta$ in, say, clockwise direction by introducing an average clockwise twist of $\Theta/L$ between neighboring layers. However, the same global twist can be achieved by a counter-clockwise twist with local angle of $(2\pi-\Theta)/L$ between consecutive layers. At low temperatures, the additional free energy for a twist in the “ wrong” direction is much larger than $k_{B}T$. Therefore, this state is exponentially suppressed. Local twist angles corresponding to higher winding numbers can also be ruled out using the same argument. However, with increasing temperature, states of both chiralities (and higher winding numbers) will be mixed, leading to an extra entropic contribution to the free energy. Importantly, $\Theta=\pi$ is a special case that allows the mixing of opposite chiralities even at $T=0$. Let us now study this idea quantitatively.
@@
Figure 2: Free energy difference $\Delta F$ vs. squared twist angle $\Theta^{2}$ for size $L=40$ and several temperatures close to $T_{c}$. The dotted line shows a fit of $\Delta F$ for $T=0.80$ to a quadratic $\Theta$ dependence in the range $0\leq\Theta^{2}\leq(2\pi/3)^{2}$.
The figure shows that $\Delta F$ follows the quadratic $\Theta$ dependence up to about $\Theta=2\pi/3$ but significant deviations are observed for larger twist angles. We attribute the deviation of $\Delta F$ from the quadratic $\Theta$ dependence to the mixing of states of opposite chirality which becomes more pronounced as $\Theta$ approaches $\pi$. To quantify this mechanism, we determine the excess free energy due to the mixing. To this end, we first fit a quadratic function to $\Delta F(\Theta)$ in the range $30\leq\Theta\leq 2\pi/$ 3 (separately for each system size). Denoting the fit function by $\Delta F_{fit}(L,T,\Theta)$, we define the excess free energy $\delta\Delta F$ at twist angle $\pi$ as the difference between $\Delta F_{fit}(L,T,\pi)$ and the $\Delta F_{MC}(L,T,\pi)$ obtained from MC,
$\delta\Delta F(L,T,\pi)=\Delta F_{fit}(L,T,\pi)-\Delta F_{MC}(L,T,\pi).$ (8)
In other words, $\Delta F_{fit}$ is the free energy cost of the twist expected if only states of one chirality contribute, whereas $\delta\Delta F$ captures the additional free energy due to chirality mixing.
Figure 3 shows the excess free energy $\delta\Delta F$ for twist angle $\pi$ as a function of temperature $T$ for various system sizes.
@@
For temperatures $T>T_{c}$, in the disordered phase, $\delta\Delta F$ decreases with increasing system size and appears to vanish in the thermodynamic limit. For temperatures $T<T_{c}$, in contrast, $\delta\Delta F$ approaches a nonzero limit for $L\to\infty$. This infinite-system limit of $\delta\Delta F$ follows the function $T\ln(2)$ as temperature goes to zero  [^1] The small deviations of $\delta\Delta F$ from $T\ln(2)$ for smaller system sizes persist to the lowest temperatures. They can be explained quantitatively by contributions from the quartic terms in $(\Theta/L)$ that arise in the expansion of the zero-temperature energy difference $\Delta U=U(\Theta)-U(0)=-JL^{2}(1-\cos(\Theta/L))$..

### v3

@@ Abstract
We study the response of a two-dimensional classical XY model to a finite (non-infinitesimal) twist of the boundary conditions. We use Monte Carlo simulations to evaluate the free energy difference between periodic and twisted-periodic boundary conditions and find deviations from the expected quadratic dependence on the twist angle. Consequently, the helicity modulus (spin-stiffness) shows a non-trivial dependence on the twist angle. We show that the deviation from the expected behavior arises because of the mixing of states with opposite chirality which leads to an additional entropy contribution in the quasi-long-range ordered phase. We give an improved prescription for the numerical evaluation of the helicity modulus, and we discuss the spontaneous breaking of the chiral symmetry for anti-periodic boundary conditions. We also discuss applications to discrete spin systems and some experimental scenarios where boundary conditions with finite twist are necessary.
I Introduction
@@
The study of magnetic materials with the help of spin models such as the Ising or Heisenberg models has resulted in tremendous progress in condensed matter physics. For clean systems with short-range interactions, surface energies are negligible, and the choice of boundary conditions does not affect the thermodynamic bulk behavior. On the other hand, in disordered systems such as spin glasses, effects of various boundary conditions have been a topic of active interest [1, 2, 3]. In the presence of long-range interactions in an Ising spin system, different choices of boundary condition lead to seemingly different thermodynamic behaviors. By choosing artificial coupling-dependent boundary condition, an uncountable number of exotic spin states of the ground state can be generated at any temperature, whereas free boundary conditions, which are considered physical, do not generate the same effect [4]. In XY or Heisenberg spin glasses, small rotations at the boundary are expected to yield non-smooth changes in the ground state [5]. In these systems, different boundary conditions are advantageous for revealing the numerous phases and their physical properties [6, 7, 8].
Even in system in which the boundary conditions do not affect the bulk behavior, they can have interesting and observable consequences. For example, thermodynamic Casimir forces [9], which arise due to the confinement of critical order parameter fluctuations, were found to be affected by the boundary conditions [10, 11, 12, 13, 14]. Moreover, the response of a system to a change in boundary conditions can be employed as a tool to measure (bulk) equilibrium properties.
@@
In this paper, we reconsider the seemingly simple but surprisingly complex effects of a finite (non-infinitesimal) twist of the boundary conditions on a ferromagnetic XY system. Arguments put forward by Brown and Ciftan [15] in the context of the three-dimensional Heisenberg model suggest that the mixing of states with different chirality plays an important role and affects the observed helicity modulus. However, a quantitative understanding of this effect and its origins has not been achieved yet. We therefore study this question in detail for the two-dimensional classical XY model by means of large-scale Monte Carlo simulations. We find that the free energy cost of a non-infinitesimal twist in the boundary condition in the quasi long-range ordered (QLRO) low-temperature phase deviates from the expected quadratic dependence on the twist angle. Mixing of states with opposite chirality provides an extra entropy contribution which takes the value $\ln(2)$ (in units in which $k_{B}=1$) for a $\pi$-twist. Beyond their intrinsic interest, our results potentially apply to experiments aimed at detecting the Berezinskii–Kosterlitz–Thouless (BKT) transition [16, 17]. They are also important for systems with discrete $Z(N)$ (clock) order parameter symmetry where any twist in the boundary conditions is necessarily non-infinitesimal. In addition, our findings enable novel numerical algorithms for computing the helicity modulus in simulations.
@@
Our paper is organized as follows. In section II, we introduce the XY Hamiltonian under boundary conditions with an arbitrary twist, and we define the helicity modulus. Section III contains the details of the numerical simulations. In section IV, we present our results for the dependence of the free energy on the twist angle and discuss the resulting helicity modulus. We also analyze the spontaneous breaking of the chiral symmetry that occurs for antiperiodic boundary conditions below the BKT transition. We conclude in section V.
II The Model
@@
Here, $J>0$ denotes the ferromagnetic exchange interaction (which will be set to unity in the simulations), the sum is over pairs of nearest neighbors on a $d$-dimensional hypercubic lattice, and $\mathbf{S}=(S_{x},S_{y})$ is a two-component unit vector. Equivalently, the XY spins can be represented by their phases $\phi$, defined via $S_{x}=\cos\phi,~{}S_{y}=\sin\phi$. Twisted boundary conditions can be implemented by fixing the spins at two opposite boundaries at specific orientations (phases) with a fixed angle $\Theta$ between them. Alternatively, we consider the Hamiltonian (1) with periodic boundary conditions and modify the interactions across one of the boundaries to introduce the twist. This is achieved by replacing the interaction terms across the chosen boundary by $-J\cos(\phi_{i}-\phi_{j}-\Theta)$. We call these boundary conditions twisted-periodic, and the usual periodic boundary conditions are recovered for $\Theta=0$.
@@
In a long-range ordered or quasi long-range ordered phase, a twist in the boundary conditions increases the system’s free energy. The system can lower its free energy by distributing the total twist (angular difference) $\Theta$ over the entire sample, i.e., by gradually changing the average orientation of the spins in the bulk. For a system of linear size $L$, the lowest free energy is expected when the average phase changes by $\Theta/L$ between neighboring sites (in the direction the twist is applied). For large $L$, this local phase change is small, which suggests that the free energy can be expanded in powers of $\Theta/L$. As the free energy difference $\Delta F$ between the twisted and untwisted systems must be an even function of $\Theta$, the expansion is expected to take the form
@@
$\rho_{s}(\Theta)=\frac{2\Delta F}{\Theta^{2}}L^{2-d}~{}.$ (3)
This definition still depends on the value of the imposed twist angle $\Theta$. Some papers in the literature including the seminal work by Fisher, Barber, and Jasnow [19] define the helicity modulus via a twist angle of $\pi$, i.e., via the free energy difference between periodic and antiperiodic boundary conditions (in analogy with the study of interfacial energies in Ising models). Other authors define the helicity modulus via the response to an infinitesimal twist (in the spirit of linear response theory),
$\rho_{s0}=\rho_{s}(0)=\left(\frac{\partial^{2}F}{\partial\Theta^{2}}\right)_{\Theta=0}L^{2-d}~{}.$ (4)
It has generally been assumed that the two definitions lead to the same stiffness values because the local, layer-to-layer twist $\Theta/L$ is small in either case (in the thermodynamic limit), justifying the expansion (2) of the free energy [20, 21]. However, we will see that this is not the case, at least not in two space dimensions.
The definition (4) has the advantage that the second derivative of the free energy can be expressed in terms of appropriate correlation functions of the untwisted system. This leads to the formula [22]
$\left(\frac{\partial^{2}F}{\partial\Theta^{2}}\right)_{\Theta=0}=\frac{1}{L^{2-d}}\sum_{<ij>}J\left\langle\cos(\phi_{i}-\phi_{j})\right\rangle(x_{i}-x_{j})^{2}$
$-\frac{\beta}{L^{2-d}}\left\langle\left\{\sum_{<ij>}J\sin(\phi_{i}-\phi_{j})(x_{i}-x_{j})\right\}^{2}\right\rangle$ (5)
where $x_{i}$ is the coordinate of site $i$ in the twisted direction, and delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the thermodynamic average evaluated at $\Theta=0$. This formula allows the evaluation of the helicity modulus without actually having to apply twisted boundary conditions.
As the paramagnetic phase is insensitive to the boundary conditions, the free energy difference decays exponentially with system size, $\Delta F\sim e^{-L/\xi}$ , where $\xi$ is the correlation length. This implies $\rho_{s}=0$ in the paramagnetic phase in the thermodynamic limit. In contrast, $\Delta F$ is expected to scale as $L^{d-2}$ in an ordered or quasi long-range ordered phase, and the helicity modulus is finite. In two dimensions, $\rho_{s}$ is known to have an universal jump at the BKT phase transition. For $d>2$, $\rho_{s}$ vanishes continuously at $T_{c}$, governed by critical behavior of the phase transition. In the rest of the paper, we focus on two dimensions, but we will comment on higher dimensions in the concluding section.
III Numerical Simulations
We perform large-scale Monte Carlo simulations to evaluate the free energy difference between systems with periodic and twisted-periodic boundary conditions. The free energy cannot be measured directly in a standard Monte Carlo simulation. Instead, it can be evaluated explicitly by integrating the internal energy $U=\langle H\rangle$ over the inverse temperature $\beta=1/T$,
@@ III Numerical Simulations
Our simulations employ both the single spin-flip Metropolis algorithm [25, 26] and the Wolff cluster-flip algorithm [27]. For systems with periodic boundary conditions, the efficient Wolff algorithm greatly reduces critical slowing down. Thus, a full MC sweep consists of one Metropolis sweep followed by one Wolff sweep for the case of periodic boundary conditions. However, for twisted-periodic boundary conditions with twist angle $0<\Theta<\pi$, the Wolff algorithm cannot be employed. This stems from the fact that the angle between two spins of a Wolff cluster is not preserved (but rather changes sign) when the cluster is flipped. Consequently, the cluster flip changes the energy of a twisted bond inside the cluster, invalidating the algorithm. In the case of twisted-periodic boundary conditions with $0<\Theta<\pi$, we therefore only employ Metropolis sweeps. For a twist angle of exactly $\pi$, the twisted bonds effectively become antiferromagnetic as $\cos(\Delta\phi-\pi)=-\cos(\Delta\phi)$. The energy of an antiferromagnetic bond is invariant under a sign change of $\Delta\phi$, and the Wolff algorithm can be used.
@@
For periodic boundary conditions, we perform up to $2000$ full equilibration sweeps (each consisting of a Metropolis sweep followed by a Wolff sweep) and up to $8000$ full measurement sweeps at each temperature step. For twisted-periodic boundary conditions, we perform up to about 80 0008000080\,00080 000 (Metropolis only) equilibration sweeps and up to about 160 000160000160\,000160 000 measurement sweeps. (As usual, the quality of the equilibration was confirmed by comparing the results of hot and cold starts.) We simulate systems of linear sizes up to $L^{2}=80^{2}$ and average the results over about $5000$ samples. Each sample is subjected to periodic and twisted-periodic boundary conditions, and the resulting free energies are compared to evaluate the helicity modulus from eq. (3).
@@
We use an ensemble method (see, e.g., Ref. [29]) to estimate the error of the free energy. We generate a large ensemble of synthetic internal energy curves ${U_{i}(T)=U(T)+r(T)~{}\Delta U(T)}$, where $\Delta U$ is the statistical error obtained from Monte Carlo and $r(T)$ is a random number chosen from a normal distribution of unit variance. Integrating these curves via (6) generates an ensemble of free energies $F_{i}$. Mean and standard deviation of this ensemble are then propagated through eq. (3) to find the helicity modulus and the associated error.
IV Results
IV.1 Free energy response to finite twist
@@
We now turn to our results for the response of the two-dimensional XY model to various twists in the boundary conditions. Figure 1 shows the helicity modulus $\rho_{s}(\Theta)$ as a function of temperature for three different cases, viz., an infinitesimal twist, a twist of $\Theta=\pi/6$, and a twist of $\Theta=\pi$.
@@
Figure 1: Helicity modulus $\rho_{s}(\Theta)$ as a function of temperature $T$ for different system sizes $L$. The filled symbols show $\rho_{s}$ for a twist of $\Theta=\pi/6$, obtained from the definition eq. (3) using the free energies eq. (6). The dotted lines show $\rho_{s0}$ for an infinitesimal twist measured in the untwisted system via eq. (5). The helicity modulus values for these two cases agree within their statistical errors (which are comparable to or smaller than the symbol sizes). The straight solid line corresponds to $\rho_{s}=2T/\pi$, it intersects the stiffness curves at temperatures $T^{*}(L)$. The critical temperature $T_{c}$ is obtained by extrapolating $T^{*}(L)$ to the thermodynamic limit according to eq. (7). This extrapolation gives $T_{c}=0.891(7)$. In contrast, the helicity modulus values for a twist of $\Theta=\pi$, shown as dash-dotted line for $L=40$, are significantly lower than the other data.
@@
For the infinitesimal twist, $\rho_{s}$ is measured in the untwisted system via eq. (5). For the finite twists, $\rho_{s}$ is obtained from the free energy difference between simulations with twisted and untwisted boundary conditions, as explained in Sec. III. The resulting $\rho_{s}$ values for the infinitesimal twist and $\Theta=\pi/6$ agree within their statistical errors, giving us additional confidence in our numerical approach.
@@
where $A,b$ are non-universal fitting parameters. We find $T_{c}=0.891(7)$ from twisted boundary conditions with $\Theta=\pi/6$, which agrees with high-accuracy results in the literature [28, 31, 32, 33].
We also studied a twist of $\Theta=\pi$. Unexpectedly, the helicity modulus values resulting from eq. (3) in this case are significantly below the values for smaller $\Theta$, as shown in Fig. 1. It is worth emphasizing that this happens even though the layer-to-layer twist $\Theta/L$ remains small compared to unity, justifying the expansion that leads to eq. (2). What is the reason for this surprising discrepancy? Arguments put forward in Ref. [15] suggest that the chirality of the twist plays an important role.
So far we have considered thermodynamic states that fulfill the global twist $\Theta$ in, say, clockwise direction by introducing an average clockwise twist of $\Theta/L$ between neighboring layers. However, the same global twist can be achieved by a counter-clockwise twist with local angle of $(2\pi-\Theta)/L$ between consecutive layers. At low temperatures, the additional free energy for a twist in the “wrong” direction is much larger than $k_{B}T$. Therefore, this state is exponentially suppressed. Local twist angles corresponding to higher winding numbers can also be ruled out using the same argument. However, with increasing temperature, states of both chiralities (and higher winding numbers) will be mixed, leading to an extra entropic contribution to the free energy. Importantly, $\Theta=\pi$ is a special case that allows the mixing of opposite chiralities even at $T=0$. Let us now quantitatively study the mixing of the states as twist, temperature and system size is varied.
@@
Figure 2: Free energy difference $\Delta F$ vs. squared twist angle $\Theta^{2}$ for size $L=40$ and several temperatures close to $T_{c}$. The dotted line shows a fit of $\Delta F$ for $T=0.82$ to a quadratic $\Theta$ dependence in the range $0\leq\Theta^{2}\leq(2\pi/3)^{2}$. $\Delta F$ at $T_{c}\approx 0.893$, obtained using eq. (8) is marked by the star symbol.
The figure shows that $\Delta F$ follows the quadratic $\Theta$ dependence up to about $\Theta=2\pi/3$ but significant deviations are observed for larger twist angles. The special case of anti-periodic boundary conditions ($\Theta=\pi$) right at the BKT transition temperature was already studied by Hasenbusch [34, 35]. He derived an expression for the ratio of the partition functions with periodic and antiperiodic boundary conditions in the two-dimensional XY model. It includes the leading finite-size corrections and reads
$\frac{Z_{\Theta=\pi}}{Z_{\Theta=0}}=0.08643(1)-\frac{0.1358(1)}{\ln(L)+C}~{},$ (8)
where the constant $C$ approximately captures contributions from higher order terms. For the purpose of comparing with our Monte Carlo results, we set $C=4.3$ as in Ref. [34]. The free energy difference $\Delta F$ resulting from this formula agrees well with our data, see Fig. 2.
We attribute the deviation of $\Delta F$ from the quadratic $\Theta$ dependence to the mixing of states of opposite chirality which becomes more pronounced as $\Theta$ approaches $\pi$. To quantify this mechanism, we determine the excess free energy due to the mixing. To this end, we first fit a quadratic function to $\Delta F(\Theta)$ in the range $30\leq\Theta\leq 2\pi/$ 3 (separately for each system size). Denoting the fit function by $\Delta F_{\mathrm{fit}}(L,T,\Theta)$, we define the excess free energy $\delta\Delta F$ at twist angle $\pi$ as the difference between $\Delta F_{\mathrm{fit}}(L,T,\pi)$ and the $\Delta F_{\mathrm{MC}}(L,T,\pi)$ obtained from MC,
$\delta\Delta F(L,T,\pi)=\Delta F_{\mathrm{fit}}(L,T,\pi)-\Delta F_{\mathrm{MC}}(L,T,\pi).$ (9)
In other words, $\Delta F_{\mathrm{fit}}$ is the free energy cost of the twist expected if only states of one chirality contribute, whereas $\delta\Delta F$ captures the additional free energy due to chirality mixing.
@@
Figure 3 shows the excess free energy $\delta\Delta F$ for twist angle $\pi$ as a function of temperature $T$ for various system sizes. For temperatures $T>T_{c}$, in the disordered phase, $\delta\Delta F$ decreases with increasing system size and appears to vanish in the thermodynamic limit. For temperatures $T<T_{c}$, in contrast, $\delta\Delta F$ approaches a nonzero limit for $L\to\infty$. This infinite-system limit of $\delta\Delta F$ follows the function $T\ln(2)$ as temperature goes to zero  [^1] The small deviations of $\delta\Delta F$ from $T\ln(2)$ for smaller system sizes persist to the lowest temperatures. They can be explained quantitatively by contributions from the quartic terms in $(\Theta/L)$ that arise in the expansion of the zero-temperature energy difference $\Delta U=U(\Theta)-U(0)=-JL^{2}(1-\cos(\Theta/L))$..